OpenStudy (anonymous):

would 16- 2 radical 2 be the correct answer to (4+radical 8)(4- radical 8)

OpenStudy (dape):

Nope, \[ (4+\sqrt{8})(4-\sqrt{8}) = 4^2-\sqrt{8}^2 = 16-8 = 8 \]

OpenStudy (anonymous):

alright thank you